# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v1

@@ Acknowledgements.
We are grateful to Nirmalya Kajuri, Cynthia Keeler, Zhi Li, Hervé Partouche, Sabrina Pasterski, Jan Pieter van der Schaar, and Beni Yoshida for helpful discussions. V.F. and T.M. would like to thank the Berkeley Center for Theoretical Physics for their hospitality during the early stages of this work. V.F. would like to thank Perimeter Institute for their hospitality during the final stage of this work. This research is partially supported by the Cyprus Research and Innovation Foundation grant EXCELLENCE/0421/0362. This research was also supported in part by the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research, Innovation and Science. This work was supported by JSPS KAKENHI Grant Number 23KJ1154, 24K17047.
Appendix A Formal aspects of holographic entropy in de Sitter
We review the covariant holographic entanglement prescription of Franken:2023pni in Section A.1. In Appendix A.2, we define the definitions of three alternative prescriptions to find the codimension-two surfaces computing entropies. The first one, extremality, corresponds to the direct adaptation of the Hubeny-Rangamani-Takayanagi (HRT) prescription in AdS/CFT Hubeny:2007xt . The second one, maximin, is the direct adaptation of the maximin procedure which is equivalent to the HRT prescription in AdS spacetime. The third one, C-extremality, is the proposed adaptation of extremality motivated by the non-existence of extremal surfaces in some cases in de Sitter spacetime Franken:2023pni . We prove their inequivalence and that a maximin and C-extremal surface is an extremal surface. We finally show that the three prescriptions are equivalent in the interior of the holographic screen of an observer in asymptotically de Sitter spacetime.
A.1 Holographic entanglement entropy
Inspired by the Ryu-Takayanagi (RT) formula and its generalization called the HRT formula Ryu:2006bv ; Hubeny:2007xt ; Wall:2012uf ; Faulkner:2013ana ; Engelhardt:2014gca , one can consider the problem of computing entanglement entropies in the holographic dual theory in terms of bulk quantities. This question was tackled in Susskind:dSentropy ; Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni and we review the covariant proposal of Franken:2023pni , which is used in Section 6.
@@ A.1 Holographic entanglement entropy
Figure 10: Penrose diagram of dS. The two blue lines are the two screens, and the three parts of a complete slice $\Sigma$ are depicted by black lines. The worldlines of two antipodal observers follow the two vertical lines. The bulk regions $D_{I}$ are depicted by shaded regions. For $D_{E}$ to be encoded on $\mathcal{S}_{E}$, the two screens need to be located at a Planck length away from the horizons.
The holographic screens associated with two antipodal observers are denoted as $\mathcal{S}_{L}$ and $\mathcal{S}_{R}$. Each screen is an arbitrary convex timelike surface that resides in each static patch. In this appendix, we consider global dS, including the region beyond a single static patch. Thus, we slightly generalize the notation such that $\mathcal{S}$ denotes the union of the two screens, i.e. $\mathcal{S}=\mathcal{S}_{L}\cup\mathcal{S}_{R}$. We note an arbitrary spacelike slice of $\mathcal{S}_{L}\cup\mathcal{S}_{R}$ by $\Sigma|_{\mathcal{S}_{L}}\cup\Sigma|_{\mathcal{S}_{R}}$, associated with a complete achronal slice $\Sigma$. In particular, $\Sigma|_{\mathcal{S}_{L}}$ and $\Sigma|_{\mathcal{S}_{R}}$ are spacelike regions and they are spacelike separated. We denote the boundary of the middle region (exterior) by $\mathcal{S}_{E}=\mathcal{S}_{L}\cup\mathcal{S}_{R}$. We denote $\Sigma_{I}$, with $I=L,R,E$, any achronal slice such that $\partial\Sigma_{I}=\mathcal{S}_{I}$ for $i\in\{L,E,R\}$. (Note that $\Sigma$ in the main body is now denoted by $\Sigma_{R}$.) We denote $D_{I}=D(\Sigma_{I})$ the bulk region spanned by slices $\Sigma_{I}$. Now let us consider any spacelike subregion $A$ of $\Sigma|_{\mathcal{S}_{L}}\cup\Sigma|_{\mathcal{S}_{R}}$.
Definition A.1 (Homology condition).
@@
We now review the proposal of covariant holographic entanglement entropy proposal from Franken:2023pni , based on Shaghoulian:2021cef . One of the main motivations for this conjecture comes from wedge holography Shaghoulian:2021cef ; Geng:2020fxl ; Mollabashi:2014qfa ; Miao:2020oey ; Akal:2020wfl ; Bousso:2020kmy and the consistency of its results with the Gibbons-Hawking formula as well as strong subadditivity and entanglement wedge nesting Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni .
Conjecture A.1 (Covariant holographic entanglement entropy Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni ).
@@
where $\gamma_{e}(A;D_{I})$ is the minimal extremal surface $D_{I}$-homologous to $A$ [^24] When there is no such surface in some region $D_{I}$, one has to extend the notion of extremal surface, see Appendix A.2 and Franken:2023pni . We show in Appendix A.3 that the usual notion of extremality is enough in all cases considered in this paper. The entanglement wedge $\mathcal{W}(A)$ of $A$ is given by the union of the causal diamonds of all three homology regions $\mathcal{C}_{I}(A)$. Each of them is denoted as $\mathcal{W}_{I}(A)=D(\mathcal{C}_{I}(A))$
@@
In the definitions above, we considered a restricted version of the usual HRT formula, in analogy with the restricted maximin procedure Marolf:2019bgj ; Grado-White:2020wlb , which is analogous to the original maximin definition, except that it considers bulk surfaces defined on achronal slices whose boundary is identified with the screen $\Sigma|_{\mathcal{S}}$, instead of slices with the looser condition that the boundary contains the subsystem $A$. In AdS the results agree when the maximin surface lies in a smooth region of spacetime Marolf:2019bgj . Considering restricted maximin or HRT is necessary in some backgrounds to ensure strong subadditivity and entanglement wedge nesting Grado-White:2020wlb . From another perspective, restricted extremization is important when dealing with systems that are not isolated, such as evaporating black holes where time evolution on the boundary alters the entropy Penington:2019npb ; Almheiri:2019psf . Similarly, in our scenario, each region $D_{I}$ is not isolated, motivating the use of the restricted prescription (Conjecture A.1). For instance, when considering one holographic screen, applying a unitary transformation to the complementary screen could affect the entropy through the exterior region.
@@
Conjecture A.1 is central in Section 6 when discussing the connected wedge theorem in de Sitter spacetime. As noted in Franken:2023pni , defining a covariant holographic entanglement entropy prescription in the context of doubled static patch holography leads to a number of subtleties. In particular, there are situations where the HRT-like formula does not have any solution. A solution to this problem was proposed in Franken:2023pni . We review this prescription in the following section, in addition to another possibility which is to adapt the maximin prescription of Wall:2012uf to de Sitter spacetime. There are tensions between these three alternative adaptations of the holographic entanglement entropy prescription in AdS/CFT, as studied in detail in the next section. Moreover, we prove in Appendix A.3 that they are equivalent in the static patch of an observer. This fact is essential for a proof of the connected wedge theorem.
@@
The goal of this paper is not to prove which is the correct prescription. To answer this question one should derive it from a path integral computation. We do not have any particular reason to expect that the maximin method should be equivalent to any holographic entropy prescription. On the other hand, we expect the extremization process to be involved in holographic entropy prescriptions in non-AdS spacetimes. Indeed, a great number of results in holography, such as entanglement wedge nesting, the boundary causality condition, the inclusion of the causal wedge in the entanglement wedge Akers:2016ugt , as well as the connected wedge theorem May:2019yxi ; May:2019odp , heavily rely on the fact that extremal surface is of vanishing (quantum) expansion in all directions. [^25] In particular, all these properties rely on the classical focusing theorem or restricted quantum focusing conjecture Bousso:2015mna ; Shahbazi-Moghaddam:2022hbw , which constrains the evolution of the expansion or quantum expansion along null hypersurfaces. These conditions emerge from fundamental constraints such as bulk and boundary causality, and we expect them to be satisfied beyond AdS/CFT.
A.2 Extremal, maximin, and C-extremal surfaces
@@
When the bulk region $D$ is not the whole spacetime, the extremization problem may not have a solution. This can happen in the holographic entanglement entropy prescription we are about to define Franken:2023pni . So we must relax our condition of extremality to take into account surfaces that have the smallest or largest area in $D$ without being extremal. In this case, the surface must lie at least partially on the boundary of $D$ otherwise it would be a true extremum of the area functional. Since the usual HRT prescription asks to select the minimal area surface among the extremal surfaces, we only select the surface with the smallest area in $D$.
Definition A.3.
@@
Here, we do not mean by minimal or maximal area surface in $D$ that the surface is an extremum of the area function, simply that the area takes its minimal or maximal value in $D$. If there exists a C-extremal surface satisfying the second condition of Definition A.3, it must be the minimal C-extremal surface.
@@
This definition differs from the one of Wall:2012uf by the fact the achronal slices considered here are not complete. This leads to the possibility that $\gamma_{\min}(A;\Sigma)$ is not an extremum with respect to variations along $\Sigma_{I}$.
Definition A.5.
[Maximin] Let $\Sigma|_{\mathcal{S}}$ be a spacelike slice of some holographic screen $\mathcal{S}$ and $A\in\Sigma|_{\mathcal{S}}$ some subsystem. Denote $D$ the bulk region whose spatial boundary is $\Sigma|_{\mathcal{S}}$. The maximin surface associated with $A$, denoted $\gamma_{m}(A;D)$, is the surface $\gamma_{\min}(A;\Sigma^{\prime})$ which is maximal when varying over all slices $\Sigma^{\prime}$ such that $\partial\Sigma^{\prime}=\Sigma|_{\mathcal{S}}$. The achronal slice on which $\gamma_{m}(A;D)$ is minimal is denoted $\Sigma_{m}(A;D)$. Additionally, $\gamma_{m}(A;D)$ must be stable in the sense that considering a small variation of $\Sigma_{m}(A;D)$, the new slice must have a local minimum (possibly non-extremal) in the neighborhood of $\gamma_{m}(A;D)$, with no greater area.
We will not prove the existence of maximin surfaces, as it goes well beyond the scope of this paper. Note that it was proven in some classes of asymptotically AdS spacetime Wall:2012uf ; Marolf:2019bgj . We will always assume that it exists, which does not seem to be an extravagant assumption at least in the simple examples we consider in this paper, that is connected arcs, and possibly unions of them, of spacelike slices of the horizon/stretched horizon in three-dimensional de Sitter spacetime.
A.3 Proofs of (in)equivalence
@@
Lemma A.1.
A codimension-two spacelike surface $\gamma_{\min}(A;D_{I})$ that has the smallest area in bulk region $D_{I}$ and is $D_{I}$-homologous to $A$ is the minimin surface of $A$ in region $D_{I}$, and conversely. [^27] Here minimin is defined by changing the maximization of Definition A.5 by a minimization.
@@
Proposition A.2 (Maximin+Extremality along $\Sigma_{m}$Rightarrow Extremality).
A surface $\gamma_{I}\in D_{I}$ is the extremal surface of $A$ if it is the maximin surface of $A$ and is extremal along $\Sigma_{m}(A;D_{I})$. The converse does not hold.
@@
Proof of Theorem $15$ of Wall:2012uf applies directly here when the maximin surface is extremal along $\Sigma_{m}$. The main assumptions used in Wall:2012uf are that the bulk spacetime is classical, smooth, satisfies the null energy condition, and is globally hyperbolic. All these conditions are considered to be satisfied here, see Section 2.2. The proof does not depend on the AdS geometry and the only assumption that is nontrivial here is the existence of an extremum on every slice since the slices are not global. To avoid this difficulty we impose explicitly the extremality along $\Sigma_{m}(A;D_{I})$. However, even though we assume the maximin surface exists, there might be no maximin surface that is extremal along their $\Sigma_{m}(A;D_{I})$. In such case, the converse of the proposition is trivially not satisfied, and the extremal surface is a minimax surface. ∎
An example of the converse of Proposition A.2 being false is the extremal surface in the exterior region associated with one of the screens. The extremal surface homologous to a screen in the exterior is the bifurcate horizon. However, the stable maximin surface associated with the screen is always the screen itself or the complementary screen. The bifurcate horizon is an unstable maximin surface, but a stable minimax surface.
Theorem A.3.
@@
Minimin surfaces are not maximin surfaces in general so a C-extremal surface is not always a maximin surface. Conversely, if the maximin surface is not extremal along $\Sigma_{m}$ and does not lie on $\partial D$, it cannot be C-extremal. Moreover, as we already noted maximin surfaces can fail to be extremal along $\Sigma_{m}$, so being a maximin surface does not imply extremality. Conversely, Proposition A.2 states that extremality does not imply being a maximin surface. Finally, C-extremality does not imply extremality, and conversely, by Definition A.3. ∎
Theorem A.3 implies that the C-extremization and maximin prescriptions introduced in Section A.1 are not appropriate analogs of their AdS counterpart, as they fail to be equivalent. This is not necessarily a problem in the context of the connected wedge theorem since the proof only uses the fact that the extremal surface must be a minimum on at least one slice. This is the case for all C-extremal surfaces and maximin surfaces. The only important issue is that C-extremal surfaces are not always extremal, such that we cannot always apply the focusing theorem. In the rest of this section, we show that the tension between the two prescriptions is absent in the interior of the screen associated with any observer. The analysis of the exterior region is deferred to future work.
@@
Let $\Sigma$ be an achronal slice with $\partial\Sigma=\Sigma|_{\mathcal{S}}$. Let us construct an arbitrary closed codimension-two surface $\tilde{\chi}\in\Sigma$ arbitrary close to $\Sigma|_{\mathcal{S}}$, that is every point of $\tilde{\chi}$ is infinitesimally close to $\Sigma|_{\mathcal{S}}$. Consider a null hypersurface $\mathcal{L}(\Sigma|_{\mathcal{S}})$ emanating from $\Sigma|_{\mathcal{S}}$ and directed towards the interior of $\mathcal{S}$. From Definition 4.1, $\mathcal{L}(\Sigma|_{\mathcal{S}})$ is of non-positive expansion — is a lightsheet — since it is in the interior of the screen. The lightsheets discussed here are depicted in Figure 13. Consider an arbitrary codimension-two surface $\chi\in\mathcal{L}(\Sigma|_{\mathcal{S}})$, such that that on each generator of the lightsheet $\mathcal{L}(\Sigma|_{\mathcal{S}})$, there is exactly one point of $\chi$. Construct the congruence of lightrays $\mathcal{L}(\chi)$ emanating orthogonally from $\chi$ and directed towards the interior of $\mathcal{S}$. This congruence is also of non-positive expansion from Definition 4.1. We define $\tilde{\chi}=\Sigma\cap\mathcal{L}(\chi)$, which is a closed codimension-two surface on $\Sigma$. $\mathcal{L}(\Sigma|_{\mathcal{S}})$ and $\mathcal{L}(\chi)$ being of non-positive expansion, $\tilde{\chi}$ must have a smaller area than $\Sigma|_{\mathcal{S}}$. The procedure could be followed the other way around by starting from $\tilde{\chi}$. We showed that $\Sigma|_{\mathcal{S}}$ has an area greater than or equal to that of any of its infinitesimal deformation on $\Sigma$. Hence, if it is a minimal surface on some slice, it must saturate the bound and be an extremum along this slice. ∎
Figure 13: Schematic picture of the proof of Lemma A.4. One construct an arbitrary closed surface infinitesimally close to $\Sigma|_{\mathcal{S}}$ on $\Sigma$ by constructing the lightshseet $\mathcal{L}(\Sigma|_{\mathcal{S}})$ (in gray), and then a second lightsheet $\mathcal{L}(\chi)$ (in blue) emanating from an arbitrary closed surface $\chi\in\mathcal{L}(\Sigma|_{\mathcal{S}})$ infinitesimally close to $\Sigma|_{\mathcal{S}}$. One obtain the intersection $\mathcal{L}(\chi)\cap\Sigma$ which is always smaller than $\Sigma|_{\mathcal{S}}$.
Theorem A.5.

### v3

@@ Acknowledgements.
We are grateful to Sergio Aguilar-Gutierrez, Nirma lya Kajuri, Cynthia Keeler, Josh Kirklin, Zhi Li, Alex May, Hervé Partouche, Sabrina Pasterski, Jan Pieter van der Schaar, Zixia Wei, and Beni Yoshida for helpful discussions. V.F. and T.M. would like to thank the Berkeley Center for Theoretical Physics for their hospitality during the early stages of this work. V.F. would like to thank Perimeter Institute for their hospitality during the final stage of this work. This research is partially supported by the Cyprus Research and Innovation Foundation grant EXCELLENCE/0421/0362. This research was also supported in part by the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research, Innovation and Science. This work was supported by JSPS KAKENHI Grant Number 23KJ1154, 24K17047.
Appendix A Summary of notations
This paper involves several notations.
  • Plain capital letters $B;D,J$,… denote bulk subregions.
  • The lower-case Greek letter $\gamma$ denotes an achronal codimension-two surface.
  • Lower case letters $c,r,p$,… denote points in the bulk.
  • Calligraphic capital letters $\mathscr{A,L,N,R,S}$,… denote subsystems of the null membrane, an object used in the proof of Theorem 6.1.
  • $X_{M}$ with $M=0,1,2,3$ denotes the embedding coordinates of dS3.
    $\theta=\frac{1}{\mathcal{A}}\frac{d\mathcal{A}}{d\lambda},$ (56)
  • $\mathcal{I}^{\pm}$ are the past ()(-)( - ) and future (+)(+)( + ) null infinities of $dS$. Points or regions defined on $\mathcal{I}^{\pm}$ are tilded with ${}\tilde{~}{}$.
Appendix B Formal aspects of holographic entropy in de Sitter
We review the covariant holographic entanglement prescription of Franken:2023pni in Section B.1. In Appendix B.2, we define the definitions of three alternative prescriptions to find the codimension-two surfaces computing entropies. The first one, extremality, corresponds to the direct adaptation of the Hubeny-Rangamani-Takayanagi (HRT) prescription in AdS/CFT Hubeny:2007xt . The second one, maximin, is the direct adaptation of the maximin procedure which is equivalent to the HRT prescription in AdS spacetime. The third one, C-extremality, is the proposed adaptation of extremality motivated by the non-existence of extremal surfaces in some cases in de Sitter spacetime Franken:2023pni . We prove their inequivalence and that a maximin and C-extremal surface is an extremal surface. We finally show that the three prescriptions are equivalent in the interior of the holographic screen of an observer in asymptotically de Sitter spacetime.
B.1 Holographic entanglement entropy
Inspired by the Ryu-Takayanagi (RT) formula and its generalization called the HRT formula Ryu:2006bv ; Hubeny:2007xt ; Wall:2012uf ; Faulkner:2013ana ; Engelhardt:2014gca , one can consider the problem of computing entanglement entropies in the holographic dual theory in terms of bulk quantities. This question was tackled in Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni and we review the covariant proposal of Franken:2023pni , which is used in Section 6.
@@ B.1 Holographic entanglement entropy
Figure 12: Penrose diagram of dS. The two blue lines are the two screens, and the three parts of a complete slice $\Sigma$ are depicted by black lines. The worldlines of two antipodal observers follow the two vertical lines. The bulk regions $D_{I}$ are depicted by shaded regions. For $D_{E}$ to be encoded on $\mathcal{S}_{E}$, the two screens should be located at a Planck length away from the horizons.
The holographic screens associated with two antipodal observers are denoted as $\mathcal{S}_{L}$ and $\mathcal{S}_{R}$. Each screen is an arbitrary convex timelike surface that resides in each static patch. In this appendix, we consider global dS, including the region beyond a single static patch. Thus, we slightly generalize the notation such that $\mathcal{S}$ denotes the union of the two screens, i.e. $\mathcal{S}=\mathcal{S}_{L}\cup\mathcal{S}_{R}$. We note an arbitrary spacelike slice of $\mathcal{S}_{L}\cup\mathcal{S}_{R}$ by $\Sigma|_{\mathcal{S}_{L}}\cup\Sigma|_{\mathcal{S}_{R}}$, associated with a complete achronal slice $\Sigma$. In particular, $\Sigma|_{\mathcal{S}_{L}}$ and $\Sigma|_{\mathcal{S}_{R}}$ are spacelike regions and they are spacelike separated. We denote the boundary of the middle region (exterior) by $\mathcal{S}_{E}=\mathcal{S}_{L}\cup\mathcal{S}_{R}$. We denote $\Sigma_{I}$, with $I=L,R,E$, any achronal slice such that $\partial\Sigma_{I}=\Sigma|_{\mathcal{S}_{I}}$ for $i\in\{L,E,R\}$. (Note that $\Sigma$ in the main body is now denoted by $\Sigma_{R}$.) We denote $D_{I}=D(\Sigma_{I})$ the bulk region spanned by slices $\Sigma_{I}$. Now let us consider any spacelike subregion $A$ of $\Sigma|_{\mathcal{S}_{L}}\cup\Sigma|_{\mathcal{S}_{R}}$.
Definition B.1 (Homology condition).
@@
We now review the covariant holographic entanglement entropy proposal from Franken:2023pni , based on Shaghoulian:2021cef . One of the main motivations for this conjecture comes from wedge holography Shaghoulian:2021cef ; Geng:2020fxl ; Mollabashi:2014qfa ; Miao:2020oey ; Akal:2020wfl ; Bousso:2020kmy and the consistency of its results with the Gibbons-Hawking formula as well as strong subadditivity and entanglement wedge nesting Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni .
Conjecture B.1 (Covariant holographic entanglement entropy Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni ).
@@
where $\gamma_{e}(A;D_{I})$ is the minimal extremal surface $D_{I}$-homologous to $A$ [^32] When there is no such surface in some region $D_{I}$, one has to extend the notion of extremal surface, see Appendix B.2 and Franken:2023pni . We show in Appendix B.3 that the usual notion of extremality is enough in all cases considered in this paper. The entanglement wedge $\mathcal{W}(A)$ of $A$ is given by the union of the causal diamonds of all three homology regions $\mathcal{C}_{I}(A)$. Each of them is denoted as $\mathcal{W}_{I}(A)=D(\mathcal{C}_{I}(A))$
@@
In the definitions above, we considered a restricted version of the usual HRT formula, in analogy with the restricted maximin procedure Marolf:2019bgj ; Grado-White:2020wlb , which is analogous to the original maximin definition, except that it considers bulk surfaces defined on achronal slices whose boundary is identified with the screen $\Sigma|_{\mathcal{S}}$, instead of slices with the looser condition that the boundary contains the subsystem $A$. In AdS the results agree when the maximin surface lies in a smooth region of spacetime Marolf:2019bgj . The restricted maximin or HRT is necessary in some backgrounds to ensure strong subadditivity and entanglement wedge nesting Grado-White:2020wlb . From another perspective, restricted extremization is important when dealing with systems that are not isolated, such as evaporating black holes where time evolution on the boundary alters the entropy Penington:2019npb ; Almheiri:2019psf . Similarly, in our scenario, each region $D_{I}$ is not isolated, motivating the use of the restricted prescription (Conjecture B.1). For instance, when considering one holographic screen, applying a unitary transformation to the complementary screen could affect the entropy through the exterior region.
@@
Conjecture B.1 is central in Section 6 when discussing the connected wedge theorem in de Sitter spacetime. As noted in Franken:2023pni , defining a covariant holographic entanglement entropy prescription in the context of doubled static patch holography leads to several subtleties. In particular, there are situations where the HRT-like formula does not have any solution. A solution to this problem was proposed in Franken:2023pni . [^33] See Hao:2024nhd for related issues in islands computations. We review this prescription in the following section, in addition to another possibility which is to adapt the maximin prescription of Wall:2012uf to de Sitter spacetime. There are tensions between these three alternative adaptations of the holographic entanglement entropy prescription in AdS/CFT, as studied in detail in the next section. Moreover, we prove in Appendix B.3 that they are equivalent in the static patch of an observer. This fact is essential for a proof of the connected wedge theorem.
@@
The goal of this paper is not to prove which is the correct prescription. To answer this question one should derive it from a path integral computation. We do not have any particular reason to expect that the maximin method should be equivalent to any holographic entropy prescription. On the other hand, we expect holographic entropy prescriptions in non-AdS spacetimes to involve an extremization process. Indeed, a great number of results in holography, such as entanglement wedge nesting, the boundary causality condition, the inclusion of the causal wedge in the entanglement wedge Akers:2016ugt , as well as the connected wedge theorem May:2019yxi ; May:2019odp , heavily rely on the fact that extremal surface is of vanishing (quantum) expansion in all directions. [^34] In particular, all these properties rely on the classical focusing theorem or restricted quantum focusing conjecture Bousso:2015mna ; Shahbazi-Moghaddam:2022hbw , which constrains the evolution of the expansion or quantum expansion along null hypersurfaces. These conditions emerge from fundamental constraints such as bulk and boundary causality, and we expect them to be satisfied beyond AdS/CFT.
B.2 Extremal, maximin, and C-extremal surfaces
@@
When the bulk region $D$ is not the whole spacetime, the extremization problem may not have a solution. This can happen in the holographic entanglement entropy prescription we are about to define Franken:2023pni . So we must relax our condition of extremality to take into account surfaces that have the smallest or largest area in $D$ without being extremal. In this case, the surface must lie at least partially on the boundary of $D$ otherwise it would be a true extremum of the area functional. Since the usual HRT prescription asks to select the minimal area surface among the extremal surfaces, we only select the surface with the smallest area in $D$. In Hao:2024nhd , a similar formula was derived in the context of islands computations in de Sitter space.
Definition B.3.
@@
Here, we do not mean by minimal or maximal area surface in $D$ that the surface is an extremum of the area function, simply that the area takes its minimal or maximal value in $D$. If there exists a C-extremal surface satisfying the second condition of Definition B.3, it must be the minimal C-extremal surface.
@@
This definition differs from that of Wall:2012uf by the fact the achronal slices considered here are not complete. This leads to the possibility that $\gamma_{\min}(A;\Sigma)$ is not an extremum with respect to variations along $\Sigma_{I}$.
Definition B.5.
[Maximin] Let $\Sigma|_{\mathcal{S}}$ be a spacelike slice of some holographic screen $\mathcal{S}$ and $A\in\Sigma|_{\mathcal{S}}$ some subsystem. Let $D$ denote the bulk region whose spatial boundary is $\Sigma|_{\mathcal{S}}$. The maximin surface associated with $A$, denoted $\gamma_{m}(A;D)$, is the surface $\gamma_{\min}(A;\Sigma^{\prime})$ which is maximal when varying over all slices $\Sigma^{\prime}$ such that $\partial\Sigma^{\prime}=\Sigma|_{\mathcal{S}}$. The achronal slice on which $\gamma_{m}(A;D)$ is minimal is denoted $\Sigma_{m}(A;D)$. Additionally, $\gamma_{m}(A;D)$ must be stable in the sense that considering a small variation of $\Sigma_{m}(A;D)$, the new slice must have a local minimum (possibly non-extremal) in the neighborhood of $\gamma_{m}(A;D)$, with no greater area.
We will not prove the existence of maximin surfaces, as it goes well beyond the scope of this paper. Note that this was proven for certain classes of asymptotically AdS spacetime Wall:2012uf ; Marolf:2019bgj . We will always assume that it exists, which does not seem to be an extravagant assumption at least in the simple examples we consider in this paper, that is connected arcs, and possibly unions of them, of spacelike slices of the horizon/stretched horizon in three-dimensional de Sitter spacetime.
B.3 Proofs of (in)equivalence
@@
Lemma B.1.
A codimension-two spacelike surface $\gamma_{\min}(A;D_{I})$ that has the smallest area in bulk region $D_{I}$ and is $D_{I}$-homologous to $A$ is the minimin surface of $A$ in region $D_{I}$, and conversely. [^36] Here minimin is defined by changing the maximization of Definition B.5 by a minimization.
@@
Proposition B.2 (Maximin+Extremality along $\Sigma_{m}$Rightarrow Extremality).
A surface $\gamma_{I}\in D_{I}$ is the extremal surface of $A$ if it is the maximin surface of $A$ and is extremal along $\Sigma_{m}(A;D_{I})$. The converse also holds only if there exists such a maximin surface extremal along $\Sigma_{m}$.
@@
Proof of Theorem $15$ of Wall:2012uf applies directly here when the maximin surface is extremal along $\Sigma_{m}$. The main assumptions in Wall:2012uf are that the bulk spacetime is classical, smooth, satisfies the null energy condition, and is globally hyperbolic. All these conditions are considered to be satisfied here, see Section 2.1. The proof does not depend on the AdS geometry and the only assumption that is nontrivial here is the existence of an extremum on every slice since the slices are not global. To avoid this difficulty we impose explicitly the extremality along $\Sigma_{m}(A;D_{I})$. However, even though we assume the maximin surface exists, there might be no maximin surface that is extremal along their $\Sigma_{m}(A;D_{I})$. In such cases, the converse of the proposition is trivially not satisfied, and the extremal surface is a minimax surface. ∎
An example of the converse of Proposition B.2 being false is the extremal surface in the exterior region associated with one of the screens. The extremal surface homologous to a screen in the exterior is the bifurcate horizon. However, the stable maximin surface associated with the screen is always the screen itself or the complementary screen. The bifurcate horizon is an unstable maximin surface, but a stable minimax surface.
Theorem B.3.
@@
Minimin surfaces are not maximin surfaces in general so a C-extremal surface is not always a maximin surface. Conversely, if the maximin surface is not extremal along $\Sigma_{m}$ and does not lie on $\partial D$, it cannot be C-extremal. Moreover, as we already noted maximin surfaces can fail to be extremal along $\Sigma_{m}$, so being a maximin surface does not imply extremality. Conversely, Proposition B.2 states that extremality does not imply being a maximin surface. Finally, C-extremality does not imply extremality, and conversely, by Definition B.3. ∎
Theorem B.3 implies that the C-extremization and maximin prescriptions introduced in Section B.1 are not appropriate analogs of their AdS counterpart, as they fail to be equivalent. This is not necessarily a problem in the context of the connected wedge theorem since the proof only uses the fact that the extremal surface must be a minimum on at least one slice. This is the case for all C-extremal surfaces and maximin surfaces. The only important issue is that C-extremal surfaces are not always extremal, such that we cannot always apply the focusing theorem. In the rest of this section, we show that the tension between the two prescriptions is absent in the interior of the screen associated with any observer. The analysis of the exterior region is deferred to future work.
@@
Let $\Sigma$ be an achronal slice with $\partial\Sigma=\Sigma|_{\mathcal{S}}$. Let us construct an arbitrary closed codimension-two surface $\tilde{\chi}\in\Sigma$ arbitrary close to $\Sigma|_{\mathcal{S}}$, that is every point of $\tilde{\chi}$ is infinitesimally close to $\Sigma|_{\mathcal{S}}$. Consider a null hypersurface $\mathcal{L}(\Sigma|_{\mathcal{S}})$ emanating from $\Sigma|_{\mathcal{S}}$ and directed towards the interior of $\mathcal{S}$. From Definition 4.2, $\mathcal{L}(\Sigma|_{\mathcal{S}})$ is of non-positive expansion — is a lightsheet — since it is in the interior of the screen. The lightsheets discussed here are depicted in Figure 15. Consider an arbitrary codimension-two surface $\chi\in\mathcal{L}(\Sigma|_{\mathcal{S}})$, such that on each generator of the lightsheet $\mathcal{L}(\Sigma|_{\mathcal{S}})$, there is exactly one point of $\chi$. Construct the congruence of lightrays $\mathcal{L}(\chi)$ emanating orthogonally from $\chi$ and directed towards the interior of $\mathcal{S}$. This congruence is also of non-positive expansion from Definition 4.2. We define $\tilde{\chi}=\Sigma\cap\mathcal{L}(\chi)$, which is a closed codimension-two surface on $\Sigma$. $\mathcal{L}(\Sigma|_{\mathcal{S}})$ and $\mathcal{L}(\chi)$ being of non-positive expansion, $\tilde{\chi}$ must have a smaller area than $\Sigma|_{\mathcal{S}}$. The procedure could be followed the other way around by starting from $\tilde{\chi}$. We showed that $\Sigma|_{\mathcal{S}}$ has an area greater than or equal to that of any of its infinitesimal deformation on $\Sigma$. Hence, if it is a minimal surface on some slice, it must saturate the bound and be an extremum along this slice. ∎
Figure 15: Schematic picture of the proof of Lemma B.4. One constructs an arbitrary closed surface infinitesimally close to $\Sigma|_{\mathcal{S}}$ on $\Sigma$ by constructing the lightsheet $\mathcal{L}(\Sigma|_{\mathcal{S}})$ (in gray), and then a second lightsheet $\mathcal{L}(\chi)$ (in blue) emanating from an arbitrary closed surface $\chi\in\mathcal{L}(\Sigma|_{\mathcal{S}})$ infinitesimally close to $\Sigma|_{\mathcal{S}}$. The intersection $\tilde{\chi}=\mathcal{L}(\chi)\cap\Sigma$ is always smaller than $\Sigma|_{\mathcal{S}}$.
Theorem B.5.